# Matters arising: Tunnelling photons pose no challenge to Bohmian machanics

### v1

@@
Matters arising: Tunnelling photons pose no challenge to Bohmian machanics
Yun-Fei Wang    Xiao-Yu Wang    Hui Wang Hefei National Research Center for Physical Sciences at the Microscale and School of Physical Sciences, University of Science and Technology of China, Hefei 230026, China Shanghai Research Center for Quantum Science and CAS Center for Excellence in Quantum Information and Quantum Physics, University of Science and Technology of China, Shanghai 201315, China Hefei National Laboratory, University of Science and Technology of China, Hefei 230088, China    Chao-Yang Lu Hefei National Research Center for Physical Sciences at the Microscale and School of Physical Sciences, University of Science and Technology of China, Hefei 230026, China Shanghai Research Center for Quantum Science and CAS Center for Excellence in Quantum Information and Quantum Physics, University of Science and Technology of China, Shanghai 201315, China Hefei National Laboratory, University of Science and Technology of China, Hefei 230088, China New Cornerstone Science Laboratory, Shenzhen 518054, China
(July 27, 2025)
@@
where $\rho_{a}=|\psi_{a}|^{2}/(|\psi_{a}|^{2}+|\psi_{m}|^{2})$. Thus $v$ can be determined by recording the populations $|\psi_{a}|^{2}$ and $|\psi_{m}|^{2}$ in the auxiliary and main waveguides at different $x$ positions, respectively. Noting that this is not velocity measurement since the direction of motion can not be determined. The authors say that this speed “represents the local speed of particles directly behind the potential step”, but the direction of this speed is still unclear. The second velocity is determined using guiding equation $\vec{v}_{s}=\frac{\hbar}{m}\nabla S$ of BM through “interfere quantum states in the step potential with their spatially mirrored images to detect phase gradients”. Based on the authors’ claim of $v_{s}=0$ for negative kinetic energy case, we believe that this $v_{s}$ should along the $x$ direction (see Methods).
Figure 2: Quantum tunnelling between two coupled waveguides based on Bohmian mechanics. Four group of velocities at $x$=5 $\mu$m (a), 10 $\mu$m (b), 20 $\mu$m (c) and 40 $\mu$m (d) are displayed. If $\Delta<-\hbar J_{0}$, both $v_{m}$ and $v_{a}$ along $x$ are zero. In the regime of $|\Delta|<\hbar J_{0}$, $v_{m}$ and $v_{a}$ are distinguished because of the different quantum potential. For $\Delta>\hbar J_{0}$, $v_{m}$ and $v_{a}$ are identical.
@@
The authors in Ref. [3] model $\rho_{a}$ as $\rho_{a}=(J_{0}x/v)^{2}$. Though they emphasize this speed $v$ “represent the the local speed of particles directly behind the potential step”, we still have concerns whether this is acceptable. Even if accepted, $v$ cannot be directly compared to the Bohmian velocity ($v_{s}$) defined by the guiding equation. First, the direction of $v$ is ambiguous—does it include a $y$-component? In contrast, $v_{s}$ is strictly along the $x$-direction (see Methods). Second, $v_{s}$ inherently satisfies the continuity equation derived from the Schrödinger equation, whereas $v$ does not. Any speed rigorously defined within the Schrödinger formalism must be equivalent to $v_{s}$ to ensure consistency with probability conservation.
Next, we derive the Bohmian velocity $v_{s}$ from continuity equation: $v_{s}(x)=\frac{\hbar}{m}\text{Im}\frac{\psi^{*}\partial_{x}\psi}{\psi^{*}\psi}$. Using the solution in equation 3, $v_{s,m}(x)$ and $v_{s,a}(x)$ can be determined. Fig. 2 shows these calculated velocities at four position ($x=5,10,20,40$ $\mu$m). In the region of $\Delta<-\hbar J_{0}$, all velocities vanish. This is expected: the wavefunction decays exponentially (purely evanescent), indicating photons are stationary along $x$. Crucially, this does not preclude tunnelling—finite probabilities $|\psi_{a}|^{2}$ and $|\psi_{m}|^{2}$ persist in the waveguides, enabling tunnelling despite zero net flow.
To definitively demonstrate this experiment poses no challenge to BM, we prove BM yields identical photon populations $\rho_{a}$ to the Schrödinger solution (equation 4; full derivation in Methods). First, substituting the wavefunction in equations 15 into the Schrödinger equations yields new quantum Hamilton-Jacobi equations 16 and continuity equations 17. The continuity equations contain a tunnelling current term $j_{0}=2J_{0}R_{a}R_{m}\text{sin}(S_{a}-S_{m})$, which accounts for inter-waveguide quantum tunnelling while preserving probability conservation. Solving the reduced continuity equation $\partial_{x}(\rho_{a,B}v_{a,x})+j_{0}=0$ with BM-derived $v_{a,x}$ and $j_{0}$ gives $\rho_{a,B}$, which is presented in equation 22 in Methods. Subsequent analysis confirms exact agreement with the Schrödinger-predicted $\rho_{a}$ (equation 4). This result indicates that (i) the Bohmian velocity $v_{a,x}$ generates the same $\rho_{a}$ as standard Schrödinger equation, and (ii) both interpretations predict identical tunnelling dynamics, refuting claims in Ref. [3].

### v2

# arXiv:2507.20101v2[quant-ph] 02 Mar 2026

@@
arXiv:2507.20101v2[quant-ph] 02 Mar 2026
@@
where $\rho_{a}=|\psi_{a}|^{2}/(|\psi_{a}|^{2}+|\psi_{m}|^{2})$. Thus $v$ can be determined by recording the populations $|\psi_{a}|^{2}$ and $|\psi_{m}|^{2}$ in the auxiliary and main waveguides at different $x$ positions, respectively. Noting that this is not velocity measurement since the direction of motion can not be determined. The authors say that this speed “represents the local speed of particles directly behind the potential step”, but the direction of this speed is still unclear. The second velocity is determined using guiding equation $\vec{v}_{s}=\frac{\hbar}{m}\nabla S$ of BM through “interfere quantum states in the step potential with their spatially mirrored images to detect phase gradients”. Based on the authors’ claim of $v_{s}=0$ for negative kinetic energy case, we believe that this $v_{s}$ should along the $x$ direction (see Supplementary information).
[…]
@@
The authors in Ref. [3] model $\rho_{a}$ as $\rho_{a}=(J_{0}x/v)^{2}$. Though they emphasize this speed $v$ “represent the the local speed of particles directly behind the potential step”, we still have concerns whether this is acceptable. Even if accepted, $v$ cannot be directly compared to the Bohmian velocity ($v_{s}$) defined by the guiding equation. First, the direction of $v$ is ambiguous—does it include a $y$-component? In contrast, $v_{s}$ is strictly along the $x$-direction (see Supplementary information). Second, $v_{s}$ inherently satisfies the continuity equation derived from the Schrödinger equation, whereas $v$ does not. Any speed rigorously defined within the Schrödinger formalism must be equivalent to $v_{s}$ to ensure consistency with probability conservation.
[…]
To definitively demonstrate this experiment poses no challenge to BM, we prove BM yields identical photon populations $\rho_{a}$ to the Schrödinger solution (equation 4; full derivation in Supplementary information). First, substituting the wavefunction in equation (11) in SM into the Schrödinger equations yields new quantum Hamilton-Jacobi equations and continuity equations. The continuity equations contain a tunnelling current term $j_{0}=2J_{0}R_{a}R_{m}\text{sin}(S_{a}-S_{m})$, which accounts for inter-waveguide quantum tunnelling while preserving probability conservation. Solving the reduced continuity equation $\partial_{x}(\rho_{a,B}v_{a,x})+j_{0}=0$ with BM-derived $v_{a,x}$ and $j_{0}$ gives $\rho_{a,B}$, which is presented in equation (18) in Supplementary information. Subsequent analysis confirms exact agreement with the Schrödinger-predicted $\rho_{a}$ (equation 4). This result indicates that (i) the Bohmian velocity $v_{a,x}$ generates the same $\rho_{a}$ as standard Schrödinger equation, and (ii) both interpretations predict identical tunnelling dynamics, refuting claims in Ref. [3].
@@
Supplementary information for
Tunnelling photons pose no challenge to Bohmian machanics
Deriving the $\psi_{m}$, $\psi_{a}$ solution from Schrödinger equation
The two coupled Schrödinger equations in the main text satisfy:
$i\hbar\partial_{t}(\psi_{m}+\psi_{a})=$ $-\frac{\hbar^{2}}{2m}\frac{d^{2}(\psi_{m}+\psi_{a})}{dx^{2}}+V_{0}(\psi_{m}+\psi_{a}),$ (5)
$i\hbar\partial_{t}(\psi_{m}-\psi_{a})=$ $-\frac{\hbar^{2}}{2m}\frac{d^{2}(\psi_{m}-\psi_{a})}{dx^{2}}$
$+V_{0}(\psi_{m}-\psi_{a})-2\hbar J_{0}(\psi_{m}-\psi_{a}).$
We get two solutions:
$\psi_{m}+\psi_{a}$ $=A_{+}e^{ik_{+}x}+B_{+}e^{-ik_{+}x},$ (6)
$\psi_{m}-\psi_{a}$ $=A_{-}e^{ik_{-}x}+B_{-}e^{-k_{-}x},$
where
$k_{+}$ $=\sqrt{\frac{2m(E-V_{0})}{\hbar^{2}}}=\sqrt{\frac{2m(\Delta-\hbar J_{0})}{\hbar^{2}}},$ (7)
$k_{-}$ $=\sqrt{\frac{2m(E-V_{0}+2\hbar J_{0})}{\hbar^{2}}}=\sqrt{\frac{2m(\Delta+\hbar J_{0})}{\hbar^{2}}},$
$\Delta=E-V_{0}+\hbar J_{0}$, and $A_{i},B_{i}$ is the amplitude of each transmission mode.
Subsequently, the solution of $\psi_{m}$ and $\psi_{a}$ is:
$\psi_{m}$ $=\frac{1}{2}(A_{+}e^{ik_{+}x}+A_{-}e^{ik_{-}x}+B_{+}e^{-ik_{+}x}+B_{-}e^{-ik_{-}x}),$ (8)
$\psi_{a}$ $=\frac{1}{2}(A_{+}e^{ik_{+}x}-A_{-}e^{ik_{-}x}+B_{+}e^{-ik_{+}x}-B_{-}e^{-ik_{-}x}).$
According to the initial conditions that the initial auxiliary mode is zero and there is no back propagation mode, we have the following:
$A_{+}$ $=c_{0},$ (9)
$A_{-}$ $=c_{0},$
$B_{+}$ $=0,$
$B_{-}$ $=0.$
Finally two transmitted wavefunctions are given by:
$\psi_{m}$ $=\frac{c_{0}}{2}(e^{ik_{+}x}+e^{ik_{-}x}),$ (10)
$\psi_{a}$ $=\frac{c_{0}}{2}(e^{ik_{+}x}-e^{ik_{-}x}).$
This result indicates that two modes ($k_{+}$ and $k_{-}$) exist in two coupled waveguides with different energies.
Analyzing the behavior of $\rho_{a}$ when $x$ is small
Based on equation 10, the population of auxiliary mode $\rho_{a}=|\psi_{a}|^{2}$ (it can be checked that the normalization method in Ref. [3] will not change the Taylor expansion behavior at $x=0$) can be derived as:
$\rho_{a}=$ $\frac{|c_{0}|^{2}}{4}(e^{i(k_{+}-k_{+}^{*})x}-e^{i(k_{-}-k_{+}^{*})x}$ (11)
$-e^{i(k_{+}-k_{-}^{*})x}+e^{i(k_{-}-k_{-}^{*})x}).$
When $x$ is small, we can expand $\rho_{a}$ at $x=0$:
$\rho_{a}=$ $-\frac{|c_{0}|^{2}}{8}((k_{+}-k_{+}^{*})^{2}+(k_{-}-k_{+}^{*})^{2}$ (12)
$+(k_{+}-k_{-}^{*})^{2}+(k_{-}-k_{-}^{*})^{2})x^{2}+O(x^{3})$
$=$ $\frac{m|c_{0}|^{2}}{2\hbar^{2}}(|\Delta+\hbar J_{0}|-2\sqrt{\text{Max}\{\Delta^{2}-(\hbar J_{0})^{2},0\}}$
$+|\Delta-\hbar J_{0}|)x^{2}+O(x^{3}).$
Similar to Ref. [3], $\rho_{a}$ is quadratic in $x$, but the coefficient of $x^{2}$ is totally different.
When $|\Delta|>\hbar J_{0}$, this relation can be written as:
$\rho_{a}=2m|c_{0}|^{2}\frac{J_{0}^{2}}{(\sqrt{|\Delta+\hbar J_{0}|}+\sqrt{|\Delta-\hbar J_{0}|})^{2}}x^{2}$ (13)
$+O(x^{3}).$
If $|\Delta|\gg\hbar J_{0}$, our solution converges to Ref. [3]’s result. However, for $|\Delta|<\hbar J_{0}$, the expression simplifies to
$\rho_{a}=\frac{m|c_{0}|^{2}J_{0}}{\hbar}x^{2},$ (14)
which is independent of $\Delta$, producing a plateau in this regime. However, the model in Ref. [3] fails to show this plateau.
Bohmian mechanics for quantum tunnelling
Since the Schrodinger equations are not in standard form, the Bohmian mechanics equations for quantum tunnelling between two coupled waveguides should be rederived.
We start from the wavefunction in the forms of:
$\psi_{m}=R_{m}e^{iS_{m}},$ (15)
$\psi_{a}=R_{a}e^{iS_{a}},$
and substitute them into Schrodinger equation. The real parts correspond to two equations:
$-\hbar\partial_{t}S_{m}=$ $\frac{(\hbar\partial_{x}S_{m})^{2}}{2m}-\frac{\hbar^{2}}{2m}\frac{\partial_{x}^{2}R_{m}}{R_{m}}+V_{0}$ (16)
$+\hbar J_{0}\left(\frac{R_{a}}{R_{m}}\text{cos}(S_{a}-S_{m})-1\right),$
$-\hbar\partial_{t}S_{a}=$ $\frac{(\hbar\partial_{x}S_{a})^{2}}{2m}-\frac{\hbar^{2}}{2m}\frac{\partial_{x}^{2}R_{a}}{R_{a}}+V_{0}$
$+\hbar J_{0}\left(\frac{R_{m}}{R_{a}}\text{cos}(S_{m}-S_{a})-1\right),$
which are the Hamilton-Jacobi equations of the hidden variable particles. Here, $\hbar\partial_{t}S_{i}$ is the total energy of particles of these two modes, $\frac{(\hbar\partial_{x}S_{i})^{2}}{2m}$ is the kinetic energy, $\frac{\hbar^{2}}{2m}\frac{\partial_{x}^{2}R_{i}}{R_{i}}$ is the quantum potential, $V_{0}$ is the external potential, and $\hbar J_{0}(\frac{R_{j}}{R_{i}}\text{cos}(S_{j}-S_{i})-1)$ is the coupling energy between these two modes.
The imagine parts also correspond to two equations:
$\partial_{t}(R^{2}_{m})+\partial_{x}\left(\frac{R^{2}_{m}\hbar\partial_{x}S_{m}}{m}\right)+2J_{0}R_{a}R_{m}\text{sin}(S_{m}-S_{a})=0,$ (17)
$\partial_{t}(R^{2}_{a})+\partial_{x}\left(\frac{R^{2}_{a}\hbar\partial_{x}S_{a}}{m}\right)+2J_{0}R_{m}R_{a}\text{sin}(S_{a}-S_{m})=0,$
which are the continuity equations for two coupled waveguide system. Noting that (i) another term appears in continuity equations, which describe the coupling between two waveguides; and (ii) the $x$-direction velocity of these two modes can still be written as:
$v_{x,m}$ $=\frac{\hbar}{m}\partial_{x}S_{m}=\frac{\hbar}{m}\text{Im}\frac{\psi^{*}_{m}\partial_{x}\psi_{m}}{\psi^{*}_{m}\psi_{m}},$ (18)
$v_{x,a}$ $=\frac{\hbar}{m}\partial_{x}S_{a}=\frac{\hbar}{m}\text{Im}\frac{\psi^{*}_{a}\partial_{x}\psi_{a}}{\psi^{*}_{a}\psi_{a}}.$
Solving $\rho_{a,B}$ using Bohmian mechanics
Since the transmitted wave is monochromatic wave, the time evolution part is $e^{-i\omega t}$, so $\rho$ doesn’t contain $t$, namely $\partial_{t}\rho=0$. The remaining part of the continuity equation of auxiliary mode is :
$\partial_{x}(\rho_{a,B}v_{a,x})+j_{0}=0,$ (19)
where $j_{0}=2J_{0}R_{m}R_{a}\text{sin}(S_{a}-S_{m})=-iJ_{0}(\psi_{a}\psi^{*}_{m}-\psi^{*}_{a}\psi_{m})$.
When $x$ is sufficiently small, a Taylor expansion of $v_{a,x}$ and $j_{0}$ at $x=0$ leads to:
$v_{a,x}$ $=\frac{\hbar}{m}\text{Im}(i\frac{k_{-}+k_{+}}{2}+O(x^{1})),$ (20)
$j_{0}$ $=|c_{0}|^{2}(\frac{J_{0}}{2}(-k_{-}-k_{-}^{*}+k_{+}+k_{+}^{*})x+O(x^{2})).$
Substituting them into the continuity equation 19, we have:
$\partial_{x}\left[\rho_{a,B}\frac{\hbar}{m}\text{Im}\left(i\frac{k_{-}+k_{+}}{2}+O(x^{1})\right)\right]$ (21)
$+|c_{0}|^{2}\frac{J_{0}}{2}(-k_{-}-k_{-}^{*}+k_{+}+k_{+}^{*})x+O(x^{2})=0.$
After integration, we have:
$\rho_{a,B}$ $=|c_{0}|^{2}\frac{\frac{J_{0}}{4}(k_{-}+k_{-}^{*}-k_{+}-k_{+}^{*})x^{2}+O(x^{3})}{\frac{\hbar}{m}\text{Im}\left(i\frac{k_{-}+k_{+}}{2}+O(x^{1})\right)},$ (22)
$=|c_{0}|^{2}\frac{mJ_{0}}{\hbar}\frac{k_{-}+k_{-}^{*}-k_{+}-k_{+}^{*}}{k_{-}+k_{-}^{*}+k_{+}+k_{+}^{*}}x^{2}+O(x^{3}).$
When $\Delta>\hbar J_{0}$, equation 22 becomes:
$\rho_{a,B}$ $=|c_{0}|^{2}\frac{mJ_{0}}{\hbar}\frac{\sqrt{\Delta+\hbar J_{0}}-\sqrt{\Delta-\hbar J_{0}}}{\sqrt{\Delta+\hbar J_{0}}+\sqrt{\Delta-\hbar J_{0}}}x^{2}+O(x^{3})$ (23)
$=2m|c_{0}|^{2}\frac{J_{0}^{2}}{(\sqrt{\Delta+\hbar J_{0}}+\sqrt{\Delta-\hbar J_{0}})^{2}}x^{2}+O(x^{3}).$
[…]
$\rho_{a}=|c_{0}|^{2}\frac{mJ_{0}}{\hbar}x^{2}+O(x^{3}).$ (24)
[…]
$\Delta=\lim_{\epsilon\rightarrow 0}\Delta+i\epsilon,$ (25)
where $\epsilon$ is a small value. Noting this is a common technique in quantum field theory, which involves selecting a complex part approaching zero through analytic continuation of real-valued quantities. The rationale here lies in the fact that, even in the case of evanescent waves where the average value of $v_{x}$ vanishes, the velocities of individual hidden variable particles conform to a distribution centered around these averages. Furthermore, this distribution must satisfy the equilibrium condition in a perturbative sense, where the infinitesimal imagine part corresponds to an infinitesimal perturbation of the propagation modes. Consequently, the particle number distribution is still required to obey the continuity equation under this velocity-dependent framework.
From $\sqrt{x+iy}=\sqrt{\frac{\sqrt{x^{2}+y^{2}}+x}{2}}+i\text{sgn}(y)\sqrt{\frac{\sqrt{x^{2}+y^{2}}-x}{2}}$ (here we take the positive sign, corresponding to the physically reasonable value in physics, i.e., the evanescent wave solution), we have:
$\lim_{\epsilon\rightarrow 0}(\sqrt{\Delta+\hbar J_{0}+i\epsilon}+(\sqrt{\Delta+\hbar J_{0}+i\epsilon})^{*})$ (26)
$=\lim_{\epsilon\rightarrow 0}2\sqrt{\frac{|\Delta+\hbar J_{0}|(1+\frac{\epsilon^{2}}{2|\Delta+\hbar J_{0}|^{2}}+O(\epsilon^{4}))+\Delta+\hbar J_{0}}{2}}$
$=\lim_{\epsilon\rightarrow 0}\sqrt{\frac{\epsilon^{2}}{|\Delta+\hbar J_{0}|}+O(\epsilon^{4})},$
and
$\lim_{\epsilon\rightarrow 0}(\sqrt{\Delta-\hbar J_{0}+i\epsilon}+(\sqrt{\Delta-\hbar J_{0}+i\epsilon})^{*})$ (27)
$=\lim_{\epsilon\rightarrow 0}2\sqrt{\frac{|\Delta-\hbar J_{0}|(1+\frac{\epsilon^{2}}{2|\Delta-\hbar J_{0}|^{2}}+O(\epsilon^{4}))+\Delta-\hbar J_{0}}{2}}$
$=\lim_{\epsilon\rightarrow 0}\sqrt{\frac{\epsilon^{2}}{|\Delta-\hbar J_{0}|}+O(\epsilon^{4})},$
so equation 22 becomes:
$\rho_{a}$ $=|c_{0}|^{2}\lim_{\epsilon\rightarrow 0}\frac{mJ_{0}}{\hbar}\frac{\sqrt{\frac{\epsilon^{2}}{|\Delta+\hbar J_{0}|}}-\sqrt{\frac{\epsilon^{2}}{|\Delta-\hbar J_{0}|}}}{\sqrt{\frac{\epsilon^{2}}{|\Delta+\hbar J_{0}|}}+\sqrt{\frac{\epsilon^{2}}{|\Delta-\hbar J_{0}|}}}x^{2}+O(x^{3})$ (28)
$=|c_{0}|^{2}\frac{mJ_{0}}{\hbar}\frac{\sqrt{|\Delta-\hbar J_{0}|}-\sqrt{|\Delta+\hbar J_{0}|}}{\sqrt{|\Delta-\hbar J_{0}|}+\sqrt{|\Delta+\hbar J_{0}|}}x^{2}+O(x^{3})$
$=2m|c_{0}|^{2}\frac{J_{0}^{2}}{(\sqrt{|\Delta+\hbar J_{0}|}+\sqrt{|\Delta-\hbar J_{0}|})^{2}}x^{2}+O(x^{3}).$
Hereto, we see that all these results are consistent with what we derived from Schrödinger equation. There is no any distinction between Copenhagen interpretion and BM on the problem of quantum tunnelling. Any velocity defined from $\rho_{a}$ must be identical in both frameworks, and they must satisfy continuity equation.